OpenStudy (anonymous):

Determine whether the series is absolutely convergent, conditionally convergent, or divergent. Posting series now.....

OpenStudy (anonymous):

\[\sum_{n=2}^{\infty} (-3n / n+ 1)^{4n}\]

OpenStudy (anonymous):

when i do these problems I end up with a whole sheet of paper. When my friend does it, he writes down one equation and says O this is divegent, or o this is convergent ect.... What am I doing wrong and how do I do this problem?

OpenStudy (amistre64):

well, the "-" in it might make it troublesome, but thats not always the case; if the function is decreasing in value each term, then it has a better possibility of converging

OpenStudy (zarkon):

use the divergence test

OpenStudy (amistre64):

if we integrate it as a continuous function, it doesnt converge. I see that much

OpenStudy (amistre64):

the innards looks to go to 3 as n got to infinity; so 3^(inf) = inf to me

OpenStudy (amistre64):

so it diverges on that account. even tho with the negative itll tend to oscilate back and for

OpenStudy (anonymous):

i get \[\infty / \infty \in the limit\]

OpenStudy (anonymous):

sorry ment in the limit

OpenStudy (zarkon):

\[\lim_{n\to\infty}\left(\frac{-3n}{ n+ 1}\right)^{4n}\neq 0\] therefore the series diverges

OpenStudy (anonymous):

what is the limit then. Just wondering.

OpenStudy (zarkon):

you could also use the root test

OpenStudy (zarkon):

the limit is infinity

OpenStudy (amistre64):

complex infinity i believe

OpenStudy (zarkon):

you can't do that

OpenStudy (amistre64):

ack!! ... i see it

OpenStudy (anonymous):

okay then since the limit doesnt equal zero it diverges. How would I have known to use the test among so many other test.

OpenStudy (zarkon):

\[\lim_{n\to\infty}\left(\left[\frac{-3n}{ n+ 1}\right]^4\right)^{n}\] I like this more

OpenStudy (amistre64):

i dont think its about which test to use so much as you have enough tools at your disposal, use them :)

OpenStudy (zarkon):

for large n \[\left[\frac{-3n}{ n+ 1}\right]^4\approx 81\] \[81^n\] gets big ;)

OpenStudy (anonymous):

thanks guys. Ill going to try one more of these then take a breakl lol. Calc make my brain hurt lol.

OpenStudy (anonymous):

I am**